OpenStudy (anonymous):

find the equation of the line through the given pair of points in standard form using only integers (5,-3) and (-2,4)

OpenStudy (anonymous):

standard form is ax+bx+cx

OpenStudy (anonymous):

stadard form is Ax+ Bx=c

OpenStudy (anonymous):

yes, don't know why I put x on but you first find the slope. the formula is y2-y1/x2-x1

OpenStudy (anonymous):

4-(-3)/-2-5= 7/-7= -1 negative 1 is the slope

OpenStudy (anonymous):

put the equation in point slope form y-(-3)=-1(x-5)= y-(-3)=-1x+5

OpenStudy (anonymous):

write in slope intercept form by just add-3 to both sides and you get y=-1x+8 add 1x to both sides and you get the equation in standard form which is 1x+y=8

OpenStudy (anonymous):

subtract 3

OpenStudy (anonymous):

you add 3 because the 3 is a negative

OpenStudy (anonymous):

no it turns positive

OpenStudy (anonymous):

it turns positive when you add it to the other side of the equation

OpenStudy (anonymous):

y--3=y+3 hence you subtract by 3 to cancel it out in that term

OpenStudy (anonymous):

where did you get the +3 at the end of the equation?